# Boundary rigidity of systolic and Helly complexes

### v1

@@ 6. Final remarks
It will be also interesting to investigate the discrete variants Conjecture 1.4 of Burago and Ivanov [9] about simple manifolds as minimal fillings. We say that a class $\mathfrak{X}$ of simplicial (or cube) complexes satisfies the minimal filling property if for any two complexes $X$ and $Y$ from $\mathfrak{X}$ such that $X$ and $Y$ have isomorphic boundaries and $d_{G(X)}(x,y)\leq d_{G(Y)}(x,y)$ for any $x,y\in\partial X\cong\partial Y$, the inequality $|X^{0}|\leq|Y^{0}|$ holds. Namely, we replace the volume of the manifold as in [9] by the number of vertices. More generally, we say that a class $\mathfrak{X}$ of simplicial (or cube) complexes satisfies the universal minimal filling property if for any complex $X\in{\mathfrak{X}}$ and any simplicial (cube) complex $Y$ such that $X$ and $Y$ have isomorphic boundaries and $d_{G(X)}(x,y)\leq d_{G(Y)}(x,y)$ for any $x,y\in\partial X\cong\partial Y$, the inequality $|X^{0}|\leq|Y^{0}|$ holds.
We conclude the paper with the following question:

### v2

@@ 6. Final remarks
A related question is whether our Theorems 1.1 and 1.3 admit a common generalization:
@@
Is there a characterization of the class of dismantlable (or weakly modular) graphs whose clique complexes are boundary rigid?
A positive answer to this question seems nontrivial to us for two reasons. First, although in both proofs we use dismantlability, in each case the distance matrix is updated using different metric properties of the respective classes of graphs. Second, finding such a common generalization would imply the existence of a metric geometry generalizing systolic and injective (Helly) geometries, which seem to be quite apart. However, if it exists, then it will offer the possibility of a unified treatment of systolic and Helly groups.
It will be also interesting to investigate the discrete variants Conjecture 1.4 of Burago and Ivanov [9] about simple manifolds as minimal fillings. We say that a class $\mathfrak{X}$ of simplicial (or cube) complexes satisfies the minimal filling property if for any two complexes $X$ and $Y$ from $\mathfrak{X}$ such that $X$ and $Y$ have isomorphic boundaries and $d_{G(X)}(x,y)\leq d_{G(Y)}(x,y)$ for any $x,y\in\partial X\cong\partial Y$, the inequality $|X^{0}|\leq|Y^{0}|$ holds. Namely, we replace the volume of the manifold as in [9] by the number of vertices. More generally, we say that a class $\mathfrak{X}$ of simplicial (or cube) complexes satisfies the universal minimal filling property if for any complex $X\in{\mathfrak{X}}$ and any simplicial (cube) complex $Y$ such that $X$ and $Y$ have isomorphic boundaries and $d_{G(X)}(x,y)\leq d_{G(Y)}(x,y)$ for any $x,y\in\partial X\cong\partial Y$, the inequality $|X^{0}|\leq|Y^{0}|$ holds. We conclude the paper with the following question:
Question 6.6.